a spaceship travel to a strange and mysterious world in an alternate universe whose elements are different from those on earth. what is the value of the missing abundance and what is the average atomic mass of this strange element.

also make a rough sketch of the mass spectrum for this element.

isotope Mass of isotope abundance of isotope(%)
1 165.9304    33.40
2 166.9320    22.90
3 167.9324    ?
Answer #1

Given:
m1 = 165.9304 amu
A1 = 33.4 %
m2 = 166.932 amu
A2 = 22.9 %
m3 = 167.9324 amu
A3 = 100 - A1 - A2
A3 = 100 - 33.4 - 22.9
A3 = 43.7 %

use:
atomic mass = sum of (mass of isotope * abundance) / 100
= (m1*A1 + m2*A2 + m3*A3)/100
= (165.9304*33.4 + 166.932*22.9 + 167.9324*43.7)/100
= 167.0346 amu

Answer:
43.7 %
167.0346 amu